OpenStudy (anonymous):

Medals 4 grabs because I realize that I keep asking the same questions,so sorry...but "Find 3f(a-4)-2h(a) if f(x)=x^2 +3x and h(x)=2x^2-3x+5"

OpenStudy (anonymous):

lets take three steps 1) first find \[3f(a-4)\] 2) then find \[2h(x)\] 3) then subtract

OpenStudy (anonymous):

\[f(x)=x^2+3x\] \[2f(a-4)=2\left( (a-4)^2+3(a-4)\right)\] \[=2\left( (a-4)(a-4)+3a-12\right)\] \[2\left(a^2-8a+16+3a-12\right)\] \[2\left(a^2-5a+4\right)\] \[2a^2-10a+8\]

OpenStudy (anonymous):

that was the hard part. easy part is \[2h(a)=2(2a^2-3a+5)\] \[=4a^2-6a+10\]

OpenStudy (anonymous):

Its 3f...

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

\[3\left(a^2-5a+4\right)\] \[3a^2-15a+12\]

OpenStudy (anonymous):

\[2a^2-15a+12-(4a^2-6a+10)\]